# FLARE: Agile Flights for Quadrotor Cable-Suspended Payload System via Reinforcement Learning

### v1

@@ Abstract
Agile flight for the quadrotor cable-suspended payload system is a formidable challenge due to its underactuated, highly nonlinear, and hybrid dynamics. Traditional optimization-based methods often struggle with high computational costs and the complexities of cable mode transitions, limiting their real-time applicability and maneuverability exploitation. In this letter, we present FLARE, a reinforcement learning (RL) framework that directly learns agile navigation policy from high-fidelity simulation. Our method is validated across three designed challenging scenarios, notably outperforming a state-of-the-art optimization-based approach by a 3x speedup during gate traversal maneuvers. Furthermore, the learned policies achieve successful zero-shot sim-to-real transfer, demonstrating remarkable agility and safety in real-world experiments, running in real time on an onboard computer. A supplementary video is available at [video [^2] https://www.youtube.com/playlist?list=PLbEQeDMEVpqxrKderAjZ3R-KmNCdQaXvUK].
{strip}[Uncaptioned image]
Figure 1: Autonomous Agile Flights of a Quadrotor Cable-suspended Payload System via Reinforcement Learning Policy. The system demonstrates its capabilities in three challenging real-world scenarios: (a) agile waypoint passing, (b) payload targeting, and (c1, c2) one-time and sequential gate traversal.
@@ I INTRODUCTION
To solve this problem, various approaches based on optimization are proposed, aiming to further exploit the system’s maneuverability for more agile flights. These approaches converge on the idea of planning motions by solving an optimal control problem (OCP) in a model-based manner. Within these approaches, the cable swing dynamics is actively leveraged to facilitate agile flights of the suspended payload system [6], even achieving collision-free agile flights in cluttered environments [7], [8]. Despite their impressive results, these works are predicated exclusively on the taut-cable paradigm, without considering the slack mode. To address this limitation, a unified system model is introduced that incorporates complementarity constraints to describe mode transitions [9]. Consequently, some works exploit the hybrid modes of this system to generate or track agile trajectories involving automatic mode switching [10], [11]. More recently, Wang et al. achieve flights with automatic mode-switching through a systematic planning and control framework, which is remarkably validated in a gate-traversal scenario [12]. However, the corresponding high computational costs prevent real-time trajectory planning in these works [9, 10, 12].
@@
Meanwhile, RL methods have been preliminarily explored for the navigation of the suspended payload system. Pioneering RL implementations prioritize the swing minimization [16], [17], achieving safety guarantees while compromising on agility. Another work augments the RL framework with meta-learning technique, facilitating rapid mid-flight adaptation to unknown physical properties [18], achieving robust performance across varying payload weights and cable lengths. However, the application of RL methods to achieve truly agile flights of the suspended payload system still remains unexplored.
Hence, in this letter, we tackle the challenge of agile flight for the suspended payload system by formulating it within the RL framework. We employ a model-free RL method in a high-fidelity, high-performance simulation platform. The trained policy network establishes a direct state-to-control mapping, circumventing the need for separate planning and control modules. Additionally, the lightweight network enables high-frequency inference solely with onboard resources, which is challenging for computationally intensive optimization-based approaches. Overall, the main contributions of this letter are summarized as follows:
@@
    We propose an RL-based navigation framework for the agile flight of the quadrotor cable-suspended payload system, featuring reward functions tailored to the characteristics of the system.
@@
    We design three representative application scenarios and instantiate the RL problem to demonstrate the method’s promising capabilities.
@@
    We validate our proposed method and its zero-shot sim-to-real transfer capability through extensive experiments conducted in both simulation and the real world. Our approach demonstrates remarkable agility and safety with high-performance, real-time execution.
@@ II METHODOLOGY
II-A Suspended-payload system dynamics
The quadrotor suspended system exhibits hybrid dynamics characterized by two modes depending on whether the cable is taut or slack.
In the taut mode, the system dynamics is governed by:
$(m_{q}+m_{p})(\boldsymbol{\ddot{x}_{p}}-\boldsymbol{g})=\left(\boldsymbol{\rho}\cdot\mathbf{R}\boldsymbol{T}-m_{q}l(\boldsymbol{\dot{\rho}}\cdot\boldsymbol{\dot{\rho}})\right)\boldsymbol{\rho}$ (1)
$\boldsymbol{x_{q}}=\boldsymbol{x_{p}}-l\boldsymbol{\rho},~~\dot{\mathbf{R}}=\mathbf{R}\boldsymbol{\hat{\omega}}$
$m_{q}l\boldsymbol{\ddot{\rho}}=-m_{q}l(\boldsymbol{\dot{\rho}}\cdot\boldsymbol{\dot{\rho}})\boldsymbol{\rho}+\boldsymbol{\rho}\times(\boldsymbol{\rho}\times\mathbf{R}\boldsymbol{T})$
$I_{q}\boldsymbol{\dot{\omega}}=\boldsymbol{M}-\boldsymbol{\omega}\times I_{q}\boldsymbol{\omega}$
where $\mathbf{T}=[0,0,T]^{T}$ represents the thrust vector, $\mathbf{R}$ denotes the rotation matrix, $\mathbf{g}$ is the gravity vector, $l$ is the cable length, $\boldsymbol{\rho}$ denotes the cable direction unit vector from quadrotor to payload. The variables $\mathbf{x}_{q}$, $\boldsymbol{\omega}$, $m_{q}$, $\mathbf{I}_{Q},\boldsymbol{M}$ represent the postion, body angular rates, mass, the inertia matrix and the moment vector of quadrotor, and $\mathbf{x}_{p}$, $m_{p}$ represent the postion and mass of payload, respectively.
In the slack mode (zero cable tension), the quadrotor resumes to nominal quadrotor dynamics, while the payload obeys free-fall dynamics.
These complex hybrid dynamics make optimization-based methods computationally expensive. Therefore, this work employs a model-free RL approach to achieve real-time navigation.
II-B Problem formulation
@@ II-B1 Observations and Actions
where $\mathbf{o}_{\textit{general}}$ represents the general observations shared across all three scenarios, while $\mathbf{o}_{\textit{specific}}$ consists of scenario-specific observations that will be detailed in section II-C.
@@
The policy network $\pi_{\psi}$ generates normalized actions $\mathbf{a}=[\tilde{T}_{\text{cmd}},\tilde{\boldsymbol{\omega}}_{\text{cmd}}]$, which are transformed to physical commands through:
@@ II-B2 Reward Function
where ${r}_{\text{general}}$ denotes the general reward component common to all three scenarios, while ${r}_{\text{specific}}$ is scenario-specific and will be introduced in section II-C.
Specifically, the general reward component is designed as ${r}_{\textit{general}}={r}_{\textit{safe}}+{r}_{\textit{crash}}+{r}_{\textit{smooth}}$, where we difine the sub-rewards as follows.
@@
The crash reward ${r}_{\text{crash}}$ penalizes boundary violations with episodic termination strategies. Specifically, the crash reward activates exclusively when the quadrotor or the payload violates the spatially invalid workspace $\mathcal{W}$, defined as:
@@ II-C Scenario customization
II-C1 Agile waypoint passing
@@
To this end, the observation vector in the waypoint-passing scenario is defined as $\mathbf{o}^{\textit{WP}}=[\mathbf{o}_{\textit{general}},\mathbf{o}^{\textit{WP}}_{\textit{specific}}]^{T}$, where $\mathbf{o}^{\textit{WP}}_{\textit{specific}}=\mathbf{o}_{w}^{\textit{WP}}$ denotes the observation of waypoints. The waypoint observation is composed of the relative position vectors, $\Delta\mathbf{x}_{q_{k}},k\in\{1,...,W\}$, from the drone to the next $W$ waypoints. We set $W=2$ to take two future waypoints into consideration, resulting in the waypoint observation vector $\mathbf{o}_{w}^{\textit{WP}}=[\Delta\mathbf{x}_{q_{1}},\Delta\mathbf{x}_{q_{2}}]^{\top}\in\mathbb{R}^{6}$.
@@
where the target reward ${r}^{\textit{WP}}_{\text{target}}$ quantifies the spatial progress toward the next waypoint, incentivizing the system to reach the target in minimum time, a key characteristic of agile flight, defined as:
$r^{\textit{WP}}_{\text{target}}=\|\Delta\mathbf{x}_{q_{t-1}}\|^{2}-\|\Delta\mathbf{x}_{q_{t}}\|^{2}.$ (11)
II-C2 Payload targeting
@@
Similarly to scenario 1, the observation vector in payload-targeting scenario is defined as $\mathbf{o}^{\textit{PT}}=[\mathbf{o}_{\textit{general}},\mathbf{o}^{\textit{PT}}_{\textit{specific}}]^{T}$, where $\mathbf{o}^{\textit{PT}}_{\textit{specific}}=\mathbf{o}_{w}^{\textit{PT}}$ denotes the observation of waypoints for the payload. Since only one waypoint is given as the payload target, we define the waypoint observation as $\mathbf{o}_{w}^{\textit{PT}}=[\Delta\mathbf{x}_{q},\Delta\mathbf{x}_{p}]^{T}$, where $\Delta\mathbf{x}_{q}$, $\Delta\mathbf{x}_{p}$ are defined as the drone’s and the payload’s relative position to the target, respectively.
@@
where the target reward ${r}^{\textit{PT}}_{\text{target}}$ in this scenario quantifies the spatial progress of the payload, instead of the quadrotor, towards the designated waypoint, defined as:
@@ II-C3 Gate traversal
The observation vector in the gate-traversal scenario is also defined as $\mathbf{o}^{\textit{GT}}=[\mathbf{o}_{\textit{general}},\mathbf{o}^{\textit{GT}}_{\textit{specific}}]^{T}$, where $\mathbf{o}^{\textit{GT}}_{\textit{specific}}=[\mathbf{o}_{w}^{\textit{GT}},\mathbf{o}_{g}^{\textit{GT}}]^{T}$ encodes the observation of waypoints and the gate position, respectively. The waypoint observation $\mathbf{o}_{w}^{\textit{GT}}$ shares the identical definiton with $\mathbf{o}_{w}^{\textit{WP}}$. Additionally, we define the gate observation $\mathbf{o}_{g}^{\textit{GT}}$ as the drone’s relative position to the gate center $\mathbf{x}_{g}$, providing information about the gate’s position.
@@
where the target reward ${r}^{\textit{GT}}_{\text{target}}$ shares the the identical definition with ${r}^{\textit{WP}}_{\text{target}}$ in Equation 11, and the gate reward ${r}^{\textit{GT}}_{\text{gate}}$ incentivizes the drone to approach and traverse the gate with a two-stage design strategy, defined as:
@@
where $r_{guide}$ incentives the magnitude of the drone’s velocity projected onto the desired direction, which is defined as a vector from the gate position to the waypoint position.
II-D Policy training
@@
$r_{\text{bound}}$ $10$ $r_{\text{excess}}$ 3
$r_{\text{arrival}}$ 2020 $\lambda_{1}$ $1\times 10^{\text{-}4}$
@@
The complex characteristics of the system typically incur a significant time cost. To address this, we leverage Genesis [19], a state-of-the-art, GPU-vectorized simulator that provides high-performance quadrotor simulation. Within this simulator, we model the system using the Unified Robot Description Format (URDF). The cable is specifically simulated as a serial chain of interconnected rigid links, an approach similar to that in [12]. All simulations are executed with a timestep of 0.01 s. This setup allows the policies for Scenarios 1, 2, and 3 to converge after approximately 0.75, 0.90, and 1.85 million timesteps, which corresponds to only 0.45, 0.52, and 2.42 hours of training time, respectively.
Moreover, the system’s complex dynamics may also introduce unmodeled factors that result in non-negligible sim-to-real gaps, such as unpredictable payload swings induced by external air disturbances and imperfect drone hovering. Therefore, we adopt a domain randomization strategy by incorporating random minor variations in the payload’s initial deviation angle, to ensure the distribution of training data encompasses real-world conditions.
@@ III SIMULATION RESULTS AND ANALYSIS
This section presents the simulation results and further analysis of our approach. In scenario 1, we evaluate the performance of our trained policy across three tracks, utilizing a near-time-optimal policy for a single drone (without a suspended payload) on the same tracks as a reference. In scenario 2, we validate the method’s ability to navigate the payload to a given target and demonstrate that the policy effectively exploits the cable’s natural swing characteristics. Finally, in scenario 3, we benchmark the performance of our method against a state-of-the-art work [12] executing identical one-time gate traversals. All training and simulation experiments are performed on a workstation with an AMD Ryzen R9-9950 CPU and NVIDIA RTX 5090 D GPU.
@@ III-A Agile waypoint passing
We first rule out a naive approach: agile flights for the suspended payload system can be achieved simply by treating the payload as an external disturbance and relying on the inherent robustness of the RL method. To verify this, we additionally establish a reference policy only incorporating the drone dynamics, similar to the methods in [13], [15]. When applied to scenario 1, the policy trained solely on the drone’s dynamics is unable to complete the waypoint-passing task, often resulting in catastrophic failure (e.g. a crash), as illustrated in Fig. LABEL:sim_task1_a. In contrast, our policy, which accounts for the complete dynamics of the coupled system, successfully navigates all waypoints.

### v2

@@ Abstract
Agile flight for the quadrotor cable-suspended payload system is a formidable challenge due to its underactuated, highly nonlinear, and hybrid dynamics. Traditional optimization-based methods often struggle with high computational costs and the complexities of cable mode transitions, limiting their real-time applicability and maneuverability exploitation. In this letter, we present FLARE, a reinforcement learning (RL) framework that directly learns an agile navigation policy from high-fidelity simulation. Our method is validated across three designed challenging scenarios, notably outperforming a state-of-the-art optimization-based approach by a 3x speedup during gate traversal maneuvers. Furthermore, the learned policies achieve successful zero-shot sim-to-real transfer, demonstrating remarkable agility and safety in real-world experiments, running in real time on an onboard computer.
@@ I INTRODUCTION
To solve this problem, various approaches based on optimization are proposed, aiming to further exploit the system’s maneuverability for more agile flights. These approaches converge on the idea of planning motions by solving an optimal control problem (OCP) in a model-based manner. Within these approaches, the cable swing dynamics is actively leveraged to facilitate agile flights of the suspended payload system [6], even achieving collision-free agile flights in cluttered environments [7], [8]. Despite their impressive results, these works are predicated exclusively on the taut-cable paradigm, without considering the slack mode. To address this limitation, a unified system model is introduced that incorporates complementarity constraints to describe mode transitions [9]. Consequently, some works exploit the hybrid modes of this system to generate or track agile trajectories involving automatic mode switching [10], [11]. More recently, Wang et al. achieve flights with automatic mode-switching through a systematic planning and control framework, named Impactor [12], which is remarkably validated in a gate-traversal scenario. However, the corresponding high computational costs prevent real-time trajectory planning in these works [9, 10, 12].
@@
Meanwhile, RL methods have been preliminarily explored for the navigation of the suspended payload system. Pioneering RL implementations prioritize the swing minimization [16], [17], achieving safety guarantees while compromising on agility. Another work augments the RL framework with the meta-learning technique, facilitating rapid mid-flight adaptation to unknown physical properties [18], achieving robust performance across varying payload weights and cable lengths. However, the application of RL methods to achieve truly agile flights of the suspended payload system still remains unexplored.
Hence, in this letter, we tackle the challenge of agile flight for the suspended payload system by formulating it within a tailored RL framework. We employ a model-free RL method in a high-fidelity simulation platform. Crucially, the trained policy establishes a direct state-to-control mapping that inherently handles the system’s hybrid dynamics (slack/taut modes), circumventing the heavy computational burden of mode-switching optimization. The lightweight network enables high-frequency inference solely with onboard resources. Overall, the main contributions of this letter are summarized as follows:
@@
    We propose an RL framework for the agile navigation of the quadrotor cable-suspended payload system, featuring observations and reward functions tailored for successful zero-shot transfer.
@@
    We design three representative application scenarios and present, to the best of our knowledge, the first real-time replanning framework that explicitly handles the system’s hybrid dynamics (taut/slack modes).
@@
    We validate our proposed method through extensive experiments along with comprehensive analysis. The results demonstrate successful zero-shot sim-to-real transfer and real-time onboard execution, notably outperforming the state-of-the-art optimization method.
@@ II METHODOLOGY
II-A Problem formulation
@@ II-A1 Observations and Actions
where $\mathbf{o}_{\textit{general}}$ represents the general observations shared across all three scenarios, while $\mathbf{o}_{\textit{specific}}$ consists of scenario-specific observations that will be detailed in section II-B.
@@
The policy network $\pi_{\psi}$ generates normalized actions $\mathbf{a}=[\tilde{T}_{\text{cmd}},\tilde{\boldsymbol{\omega}}_{\text{cmd}}]$, corresponding to the normalized collective thrust and body rates, respectively. These are transformed to physical control inputs for the quadrotor, specifically the thrust $T_{\text{cmd}}$ and body rates $\boldsymbol{\omega}_{\text{cmd}}$ through:
@@ II-A2 Reward Function
where ${r}_{\text{general}}$ denotes the general reward component common to all three scenarios, while ${r}_{\text{specific}}$ is scenario-specific and will be introduced in section II-B.
Specifically, the general reward component is designed as ${r}_{\text{general}}={r}_{\text{safe}}+{r}_{\text{crash}}+{r}_{\text{smooth}}$, where we define the sub-rewards as follows.
@@
(a) Definition and formulation of the safety reward.
The crash reward ${r}_{\text{crash}}$ penalizes boundary violations with episodic termination strategies. Specifically, the crash reward activates exclusively when the quadrotor position $\mathbf{x}_{q}$ or the payload position $\mathbf{x}_{p}$ violates the spatially valid workspace $\mathcal{W}$, defined as:
@@ II-B Scenario customization
II-B1 Agile waypoint passing
@@
To this end, the observation vector in the waypoint-passing scenario is defined as $\mathbf{o}^{\textit{WP}}=[\mathbf{o}_{\textit{general}},\mathbf{o}^{\textit{WP}}_{\textit{specific}}]^{T}$, where $\mathbf{o}^{\textit{WP}}_{\textit{specific}}=\mathbf{o}_{w}^{\textit{WP}}$ denotes the observation of waypoints. The waypoint observation is composed of the quadrotor’s relative positions, ${\color[rgb]{0,0,0}\Delta\mathbf{x}_{q}^{k}},k\in\{1,...,W\}$, to the next $W$ waypoints. We set $W=2$ to take two future waypoints into consideration, resulting in the waypoint observation vector $\mathbf{o}_{w}^{\textit{WP}}=[{\color[rgb]{0,0,0}\Delta\mathbf{x}_{q}^{1},\Delta\mathbf{x}_{q}^{2}}]^{\top}\in\mathbb{R}^{6}$.
@@
where the target reward ${r}^{\textit{WP}}_{\text{target}}$ at time step $t$ quantifies the spatial progress toward the next waypoint, incentivizing the system to reach the target in minimum time, a key characteristic of agile flight, defined as:
$r^{\textit{WP}}_{\text{target}}=\|{\color[rgb]{0,0,0}\Delta\mathbf{x}_{q_{t-1}}^{1}}\|^{2}-\|{\color[rgb]{0,0,0}\Delta\mathbf{x}_{q_{t}}^{1}}\|^{2}.$ (10)
II-B2 Payload targeting
@@
Similarly to scenario 1, the observation vector in payload-targeting scenario is defined as $\mathbf{o}^{\textit{PT}}=[\mathbf{o}_{\textit{general}},\mathbf{o}^{\textit{PT}}_{\textit{specific}}]^{T}$, where $\mathbf{o}^{\textit{PT}}_{\textit{specific}}=\mathbf{o}_{w}^{\textit{PT}}$ denotes the observation of waypoints for the payload. Meanwhile, we consider the observation of only one single future target in scenario 2. Accordingly, we define the waypoint observation as $\mathbf{o}_{w}^{\textit{PT}}=[\Delta\mathbf{x}_{q},\Delta\mathbf{x}_{p}]^{T}$, where $\Delta\mathbf{x}_{q}$, $\Delta\mathbf{x}_{p}$ are defined as the quadrotor’s and the payload’s relative position to the target, respectively.
@@
where the target reward ${r}^{\textit{PT}}_{\text{target}}$ at time step $t$ in this scenario quantifies the spatial progress of the payload, instead of the quadrotor, towards the designated waypoint, defined as:
@@ II-B3 Gate traversal
The observation vector in the gate-traversal scenario is also defined as $\mathbf{o}^{\textit{GT}}=[\mathbf{o}_{\textit{general}},\mathbf{o}^{\textit{GT}}_{\textit{specific}}]^{T}$, where $\mathbf{o}^{\textit{GT}}_{\textit{specific}}=[\mathbf{o}_{w}^{\textit{GT}},\mathbf{o}_{g}^{\textit{GT}}]^{T}$ encodes the observation of waypoints and the gate position, respectively. The waypoint observation $\mathbf{o}_{w}^{\textit{GT}}$ shares the identical definition with $\mathbf{o}_{w}^{\textit{WP}}$. Additionally, we define the gate observation $\mathbf{o}_{g}^{\textit{GT}}$ as the relative positions of the quadrotor and payload to the gate center $\mathbf{x}_{g}$, providing information about the gate’s position.
@@
where the target reward ${r}^{\textit{GT}}_{\text{target}}$ shares the identical definition with ${r}^{\textit{WP}}_{\text{target}}$ in Equation 10, and the gate reward ${r}^{\textit{GT}}_{\text{gate}}$ incentivizes the drone to approach and traverse the gate with a two-stage design strategy, defined as:
@@
where $r_{guide}$ incentivizes the magnitude of the drone’s velocity projected onto the desired direction, which is defined as a vector from the gate position to the waypoint position.
(b) Simulation results of scenario 1.
II-C Policy training
@@
$r_{{bound}}$ $10$ $r_{{excess}}$ 3
$r_{{arrival}}$ 2020 $\lambda_{1}$ $1\times 10^{-4}$
@@
Crucially, in all three scenarios, targets are uniformly resampled within a spatial volume of $2,2]\times[-2,2]\times[0.5,1.5]$ m relative to the drone, triggered by either successful arrival or episode termination, to generate random sequences rather than fixed tracks. For Scenario 3, the gate is also spatially randomized and oriented primarily perpendicular to the path for feasibility, with random yaw perturbations ($\pm 15^{\circ}$) added to enhance robustness.
The complex characteristics of the system typically incur a significant time cost. To address this, we leverage Genesis [19], a state-of-the-art, GPU-vectorized simulator that provides high-performance quadrotor simulation. The cable is specifically simulated as a serial chain of interconnected rigid links, an approach similar to that in [12]. All simulations are executed with a timestep of 0.01 s. This setup allows the policies for Scenarios 1, 2, and 3 to converge after approximately 102.0, 122.9, and 251.9 million timesteps, which corresponds to only 0.45, 0.52, and 2.42 hours of training time, respectively. Moreover, the system’s complex dynamics may also introduce unmodeled factors that result in non-negligible sim-to-real gaps, such as unpredictable payload swings induced by external air disturbances and imperfect drone hovering. Therefore, we adopt a domain randomization strategy by incorporating random minor variations in the payload’s initial deviation angle (within $\pm 10^{\circ}$), to ensure the distribution of training data encompasses real-world conditions.
@@ III SIMULATION RESULTS AND ANALYSIS
This section presents the simulation results and further analysis of our approach. In scenario 1, we evaluate the performance of our trained policy across three tracks, utilizing a near-time-optimal policy for a single drone (without a suspended payload) on the same tracks as a reference. In scenario 2, we validate the method’s ability to navigate the payload to a given target and demonstrate that the policy effectively exploits the cable’s natural swing characteristics. Finally, in scenario 3, we benchmark the performance of our method against a state-of-the-art work [12] executing identical one-time gate traversals. All training and simulation experiments are performed on a workstation with an AMD Ryzen R9-9950 CPU and NVIDIA RTX 5090 D GPU. For clarity, unless otherwise specified, all velocity metrics (Max Vel. and Avg Vel.) reported in the following tables (Table IIIV) refer to the magnitude of the quadrotor velocity.
@@ III-A Agile waypoint passing
Agility analysis: We first rule out a naive approach: agile flights for the suspended payload system can be achieved simply by treating the payload as an external disturbance and relying on the inherent robustness of the RL method. To verify this, we additionally establish a reference policy only incorporating the drone dynamics, similar to the methods in [13], [15]. When applied to scenario 1, the policy trained solely on the drone’s dynamics is unable to complete the waypoint-passing task, often resulting in catastrophic failure (e.g. a crash), as illustrated in Fig. LABEL:sim_task1_a. In contrast, our policy, which accounts for the complete dynamics of the coupled system, successfully navigates all waypoints.
@@
Robustness analysis: We further evaluate adaptability to payload mass and cable length variations. Specifically, we compare the nominal policy against policies trained with extended domain randomization (extended DR) on these two parameters. Evaluations occur at uniformly selected test points spanning $\pm 50\%$ of nominal values, involving 100 random 8-waypoint tracks for each test point. Fig. 4 shows that success rates decline with lighter payloads or longer cables, attributed to reduced inertia or larger swing range, making the payload prone to swinging above the quadrotor and causing crashes. Moreover, while the nominal policy possesses inherent robustness, extended DR amplifies it, achieving a success rate above 80% across all mass variations and boosting performance by approximately 20% on the challenging long-cable conditions, with a slightly higher average completion time in the nominal condition.
Figure 4: Robustness analysis regarding parameter variations in scenario 1. Success rates and average completion times are compared under varying payload masses (top) and cable lengths (bottom).